# Integrability from a Single Conservation Law in Quantum Spin Chains

### v3

@@ Abstract
We prove that for quantum spin chains with finite-range interactions, the existence of a specific conservation law known as the Reshetikhin condition implies the presence of infinitely many local conserved quantities, i.e., integrability. This shows that the entire hierarchy of conservation laws associated with solutions of the Yang–Baxter equation is already encoded in the lowest nontrivial conservation law. Combined with recent rigorous results on nonintegrability, our theorem strongly restricts the possibility of partially integrable systems that admit only a finite but large number of local conserved quantities. Our work establishes a rigorous foundation for the systematic identification of new integrable models and deepens the algebraic understanding of conservation-law structures in quantum spin chains.
@@ I Introduction
Quantum integrable models have long provided a unique window into the non-perturbative behavior of quantum many-body systems, from the exact solution of one-dimensional quantum magnets [1] to the emergence of generalized hydrodynamics in far-from-equilibrium dynamics [2, 3]. Although a universally accepted definition of quantum integrability is lacking [4, 5], its essential feature is often regarded as the existence of an infinite hierarchy of local conserved quantities, which we adopt as the working definition in this paper. Such conserved quantities encode much information about the initial state and constrain the dynamics in a way that generically prevents thermalization [6], enabling long-lived memory [7] and anomalous transport phenomena [8]. In contrast, generic quantum many-body systems admit only a few trivial conserved quantities, such as total energy and particle number. Therefore, integrability is a highly non-generic and fine-tuned property, and identifying integrable systems has been one of the central problems in rigorous statistical mechanics.
@@
Traditionally, one establishes integrability by solving the system explicitly; for example, solving the Yang–Baxter equation and constructing a commuting family of local charges through the logarithmic derivative of the transfer matrix [9]. As an alternative, Reshetikhin proposed a simpler criterion [10]: the existence of a single third-order conservation law with a special form. The equivalent condition had earlier appeared in the work of Jimbo and Miwa [11], where it was formulated as the tangential star-triangle relation. Unlike approaches that rely on additional algebraic structures such as the Yang–Baxter equation or the self-duality [12], this criterion is algorithmically verifiable directly from the Hamiltonian. Since this Reshetikhin condition is a necessary condition for the $R$-matrix with difference form to solve the Yang–Baxter equation, it has long served as a guide in the search for new solvable systems [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Remarkably, in all known cases, systems satisfying the Reshetikhin condition have indeed turned out to be integrable. For this reason, it has been widely regarded as a heuristic test of integrability, as proposed by Grabowski and Mathieu [25].
@@
Despite its wide use, however, it has remained an open question whether the Reshetikhin condition alone suffices to ensure integrability, namely the existence of infinitely many commuting local conserved quantities. It has often been argued that higher-order conservation laws and their mutual commutativity require additional algebraic input not captured by the third-order condition alone [17]. In this paper, we present the first complete proof that any one‑dimensional quantum system satisfying the Reshetikhin condition necessarily gives rise to an infinite sequence of commuting local charges. Specifically, we show that the so-called boost operator [26, 27, 28] allows the iterative generation of higher‑order conservation laws. We further discuss a rigorous foundation for the empirical dichotomy of integrability: either a system possesses infinitely many local conserved quantities, or a few simple at most two-body conserved charges.
@@ II Setup and main result
with $\hat{h}^{(2)}_{i}\in\mathcal{B}_{0i}\otimes\mathcal{B}_{0,i+1}$ and $\hat{h}^{(1)}_{i}\in\mathcal{B}_{0i}$. We assume Hermiticity and translational invariance. The mathematically rigorous definitions of the formal sum in Eq. (1) and of the commutators below are given in Section III.
@@ III.2 Adjoint map and conserved quantities
The map $\operatorname{ad}_{\Phi}$ is $\mathbb{C}$-linear in $\Psi$, and $\Phi\mapsto\operatorname{ad}_{\Phi}$ is also linear. Moreover, it satisfies the following properties:
@@ VI.2 Parametrized Hamiltonians
Certain integrable models, such as the Hubbard model, have an $R$-matrix that cannot be expressed as a function of a single real variable. As a result, they do not satisfy the standard Reshetikhin condition and, moreover, they admit no boost operator [37]. Instead, they are associated with families of parametrized Hamiltonians obeying a generalized Reshetikhin condition [38]. As a counterpart to Theorem 1, we show that if a parametrized Hamiltonian satisfies this generalized condition, then it is integrable.
@@
where $\lambda$ belongs to some open subset $U\subset\mathbb{R}^{D}$. For an arbitrary smooth $v:U\to\mathbb{R}^{D}$, define
@@
If $\hat{Q}^{(3)}_{v}(\lambda)$ is a conserved quantity of $\hat{H}(\lambda)$ for every $\lambda\in U$, then $\hat{Q}^{(n)}_{v}(\lambda)$ is a translationally invariant conserved quantity of $\hat{H}(\lambda)$ for all $n\geq 2$, and the operators in $\{\hat{Q}^{(n)}_{v}(\lambda)\}_{n=2}^{\infty}$ mutually commute.
@@
The above conserved quantities depend on a particular choice of $v$, but for another function $w:U\to\mathbb{R}^{D}$, it may also happen that $\hat{Q}_{w}^{(3)}(\lambda)$ is a conserved quantity [39]. In such cases, one can similarly show that $\hat{Q}_{v}^{(n)}(\lambda)$ and $\hat{Q}_{w}^{(m)}(\lambda)$ commute. More strongly, the following holds.
@@ VII Conclusion and outlook
In this paper, we have shown that a single conservation law generated by the boost operator gives rise to an infinite hierarchy of commuting local charges, or integrability. This result resolves a long-standing conjecture on quantum spin chains and provides a rigorous foundation for a heuristic sufficient criterion for integrability based on the Reshetikhin condition, independent of any additional algebraic assumptions. It also shows that the conservation-law structure of Yang–Baxter solvable models is already encoded in the lowest nontrivial conserved quantity.
It is not known whether unconventional boosts that generate conserved quantities but do not satisfy Eq. (34) also lead to integrability. Moreover, as our analysis has addressed only local conserved charges, the relation to Yang–Baxter solvability is yet to be clarified. In particular, whether one can reconstruct an $R$-matrix directly from a single conserved quantity remains a challenging problem. This problem is, in fact, equivalent to an old conjecture known as the tangential star-triangle hypothesis [11, 16].

### v4

@@ Abstract
We prove that for translationally invariant quantum spin chains with finite-range interactions, the existence of a specific conservation law known as the Reshetikhin condition implies the presence of infinitely many local conserved quantities, i.e., integrability. This shows that the entire hierarchy of conservation laws associated with solutions of the Yang–Baxter equation is already encoded in the lowest nontrivial conservation law. Combined with recent rigorous results on nonintegrability, our theorem strongly restricts the possibility of partially integrable systems that admit only a finite but large number of local conserved quantities. Our work establishes a rigorous foundation for the systematic identification of new integrable models and deepens the algebraic understanding of conservation-law structures in quantum spin chains.
@@ I Introduction
Quantum integrable models have long provided a unique window into the non‐perturbative behavior of quantum many‐body systems, from the exact solution of one‐dimensional quantum magnets [1] to the emergence of generalized hydrodynamics in far‐from‐equilibrium dynamics [2, 3]. Although a universally accepted definition of quantum integrability is lacking [4, 5], its essential feature is often regarded as the existence of an infinite hierarchy of local conserved quantities, which we adopt as the working definition in this paper. Such conserved quantities encode much information about the initial state and constrain the dynamics in a way that generically prevents thermalization [6], enabling long-lived memory [7] and anomalous transport phenomena [8]. In contrast, generic quantum many-body systems admit only a few trivial conserved quantities, such as total energy and particle number. Therefore, integrability is a highly non-generic and fine-tuned property, and identifying integrable systems has been one of the central problems in rigorous statistical mechanics.
@@
Traditionally, one establishes integrability by solving the system explicitly; for example, solving the Yang–Baxter equation and constructing a commuting family of local charges through the logarithmic derivative of the transfer matrix [9]. As an alternative, Reshetikhin proposed a simpler criterion [10]: the existence of a single third-order conservation law with a special form. The equivalent condition had earlier appeared in the work of Jimbo and Miwa [11], where it was formulated as the tangential star–triangle relation. Unlike approaches that rely on additional algebraic structures such as the Yang–Baxter equation or the self-duality [12], this criterion is algorithmically verifiable directly from the Hamiltonian. Since this Reshetikhin condition is a necessary condition for an $R$-matrix of difference form to solve the Yang–Baxter equation, it has long served as a a guiding principle in the search for new solvable systems [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Remarkably, in all known cases, systems satisfying the Reshetikhin condition have indeed turned out to be integrable. For this reason, it has been widely regarded as a heuristic test of integrability, as proposed by Grabowski and Mathieu [25].
@@
Despite its wide use, however, it has remained an open question whether the Reshetikhin condition alone suffices to ensure integrability, namely the existence of infinitely many commuting local conserved quantities. It has often been argued that higher-order conservation laws and their mutual commutativity require additional algebraic input not captured by the third-order condition alone [17]. In this paper, we present the first complete proof that any translationally invariant one‑dimensional quantum system with finite-range interactions satisfying the Reshetikhin condition necessarily gives rise to an infinite sequence of commuting local charges. Specifically, we show that the so-called boost operator [26, 27, 28] allows the iterative generation of higher‑order conservation laws. We further discuss a rigorous foundation for the empirical dichotomy of integrability: either a system possesses infinitely many local conserved quantities, or a few simple local conserved charges.
@@ II Setup and main result
with $\hat{h}^{(2)}_{i}\in\mathcal{B}_{0i}\otimes\mathcal{B}_{0,i+1}$ and $\hat{h}^{(1)}_{i}\in\mathcal{B}_{0i}$. We assume Hermiticity and translational invariance. We note that Hamiltonians with finite-range interactions that are $3$-local or higher can also be incorporated into the present framework by grouping several neighboring sites into a single effective site. The mathematically rigorous definitions of the formal sum in Eq. (1) and of the commutators below are given in Section III.
@@ III.2 Adjoint map and conserved quantities
The map $\operatorname{ad}_{\Phi}$ is $\mathbb{C}$–linear in $\Psi$, and $\Phi\mapsto\operatorname{ad}_{\Phi}$ is also linear. Moreover, it satisfies the following properties:
@@ III.3 Boost operator and main theorem
In the above discussion, we do not assume that $\Phi$ is at most $2$-local (which is equivalent to $r(\Phi)\leq 2$); rather, it is sufficient that $\Phi$ has finite-range interactions.
@@ VI.2 Parametrized Hamiltonians
Certain integrable models, such as the Hubbard model, have an $R$-matrix that cannot be expressed as a function of a single real variable. As a result, they do not satisfy the standard Reshetikhin condition and, moreover, they admit no boost operator [37]. Instead, they are associated with families of parametrized integrable Hamiltonians obeying a generalized Reshetikhin condition [38]. As a counterpart to Theorem 1, we show that if a parametrized Hamiltonian satisfies this generalized condition, then it is integrable.
@@
where $\lambda$ belongs to some open subset $U\subset\mathbb{R}^{D}$. For a smooth function $v:U\to\mathbb{R}^{D}$, define
@@
If there exists a smooth function $v:U\to\mathbb{R}^{D}$ such that $\hat{Q}^{(3)}_{v}(\lambda)$ is a conserved quantity of $\hat{H}(\lambda)$ for every $\lambda\in U$, then $\hat{Q}^{(n)}_{v}(\lambda)$ is a translationally invariant conserved quantity of $\hat{H}(\lambda)$ for all $n\geq 2$, and the operators in $\{\hat{Q}^{(n)}_{v}(\lambda)\}_{n=2}^{\infty}$ mutually commute.
We note that setting $v\equiv 0$ reproduces Theorem 1 as a special case, and is consistent with the original Reshetikhin condition for $R$-matrices of difference form.
@@
Although the above theorem focuses on a single function $v$ satisfying the generalized Reshetikhin condition, it may also happen that $\hat{Q}_{w}^{(3)}(\lambda)$ is a conserved quantity for another function $w:U\to\mathbb{R}^{D}$ [39]. In such cases, one can similarly show that $\hat{Q}_{v}^{(n)}(\lambda)$ and $\hat{Q}_{w}^{(m)}(\lambda)$ commute. More strongly, the following holds.
@@ VII Conclusion and outlook
In this paper, we have shown that a single conservation law generated by the boost operator gives rise to an infinite hierarchy of commuting local charges, or integrability. This result resolves a long-standing conjecture on quantum spin chains and provides a rigorous foundation for a heuristic sufficient criterion for integrability based on the Reshetikhin condition, independent of any additional algebraic assumptions. It also shows that the conservation-law structure of translationally invariant Yang–Baxter solvable models with finite-range interactions is already encoded in the lowest nontrivial conserved quantity.
It is not known whether unconventional boosts that generate conserved quantities but do not satisfy Eq. (34) also lead to integrability. Moreover, as our analysis has addressed only local conserved charges, the relation to Yang–Baxter solvability is yet to be clarified. In particular, whether one can reconstruct an $R$-matrix directly from a single conserved quantity remains a challenging problem. This problem is, in fact, equivalent to an old conjecture known as the tangential star–triangle hypothesis [11, 16].